# Exponents and front fluctuations in the quenched Kardar-Parisi-Zhang universality class of one and two dimensional interfaces

### v1

@@ III.3 Correlation function
Another check of the goodness of our results also comes from Eq. (18), which can be rewritten as $C_{2}(\boldsymbol{r},t)r^{-2\alpha}=g[\boldsymbol{r}/\xi(t)]$. Consequently, a plot of $C_{2}(\boldsymbol{r},t)r^{-2\alpha}$ vs. $r/\xi(t)$ should yield the universal (i.e., time-independent) scaling function $g(u)$ with asymptotic behaviors as indicated in Eq. (18). Figure 6(a) shows such a data collapse for $d=1$, for the selected times and different lattice sizes, in a doubly-logarithmic plot, in good agreement with the theoretical expectations. Note further that the $C_{2}(\boldsymbol{r},t)$ curves for different times share the same slope at small arguments; as a consequence, there is a range of $r$ where these curves partially overlap for different times. This indicates that the same scaling ansatz holds throughout the lattice, irrespective of the value of $r/\xi(t)$ and, in turn, that a single roughness exponent $\alpha$ suffices to describe the spatial properties of correlations in the system, as indeed expected within the FV dynamic scaling ansatz [Barabasi95, Krug97, Barreales20] (see also Fig. 10(a) in Appendix B).
@@
Following the same methodology used in Sec. III.1, for our simulations in $d=2$ we obtain $\theta=0.803(1)$ and $F_{c}=1.703(1)$ after fitting the $L=1024$ data to Eq. (12), see Fig. 1(b). We are not aware of previous reports in the literature of the value of $F_{c}$ in $d=2$ for the model we are presently studying.
@@
No evidence of anomalous scaling has been observed, such as, for example, the characteristic anomalous upward shift of the $C_{2}(\boldsymbol{r},t)$ curves for increasing times [Lopez97, Ramasco00, Cuerno04]. This underscores the consistency of the FV dynamic scaling ansatz for the 2D model as well and contrasts with the non-FV, anomalous scaling observed for other systems with absorbing states, see, e.g., Ref. [Dickman2000, Barreales2023] and others therein.
@@ V Statistics of fronts fluctuations
Besides the scaling exponents, the shape of the PDF for fluctuations of the front, defined according to Eq. (20) above, contributes substantially to the unambiguous identification of non-equilibrium universality classes, as has been demonstrated in several contexts [Kriecherbauer10, Halpin15, Takeuchi18, Carrasco2016, Barreales20, Rodriguez21, Marcos22, Barreales2023]. In particular, the skewness and excess kurtosis have been recently computed in a system closely related to the one we are addressing, namely, the quenched version of the Edwards-Wilkinson equation [Toivonen2022, Sillanpaa2025].
Figure 7(a) shows the PDF for the front fluctuations computed for the 1D model at different times within the growth regime, at fixed $L=10^{6}$. This plot evidences that, first, once suitably rescaled, the distribution of fluctuations indeed becomes asymptotically independent of time and system size. Second, a markedly asymmetric shape is observed, with a sharp edge for negative $\chi$ and a more regular tail for positive $\chi$. As a reference, we have plotted in Fig. 7 the Tracy-Widom distributions for the largest eigenvalue of Hermitian random matrices in the Gaussian orthogonal (TW-GOE) and unitary (TW-GUE) ensembles, as well as the familiar Gaussian distribution. The first two are known [Kriecherbauer10, Halpin15, Takeuchi18] to describe the fluctuation PDF in the standard 1D KPZ universality class, in band geometry subject to PBC and for an overall circular symmetry of the front, respectively, whereas the Gaussian distribution holds for linear models of kinetically rough surfaces [Barabasi95, Carrasco19], all of them in the case of time-dependent (rather than quenched) noise.
@@
Figure 7: PDF of the front fluctuations for the 1D model. (a) PDF computed for $L=10^{6}$ at two different times within the growth regime, see legend. The dotted lines correspond to the best fits of the tails of the PDF to Eq. (21), with the values for the tail exponents $\eta_{\pm}$ given in Table 3. (b) PDF computed at $t=50000$ for different system sizes, see legend. For both panels, the solid lines show the TW-GOE, TW-GUE, and Gaussian distributions as references.
Figure 8: PDF of the front fluctuations for the 2D model. (a) PDF computed for $L=512$ at two different times within the growth regime, see legend. The dotted lines correspond to the best fits of the tails of the PDF to Eq. (21), with the values for the tail exponents $\eta_{\pm}$ given in Table 3. (b) PDF computed at $t=600$ for different system sizes, see legend. In both panels, the solid lines show the Gaussian distribution as reference.
@@
Similar qualitative features are found for $d=2$, as shown in Fig. 8, where the PDFs exhibit quite different shapes, however. This is evidenced in Fig. 9, where $P(\chi)$ is plotted as a function of $\chi$ for both dimensionalities, with the data being available at the Zenodo open access repository [datasetchi]. Note, for the 2D KPZ universality class the PDF is known not to be in the TW family, despite being strongly non-Gaussian [Halpin-Healy2012, Halpin-Healy2013, Oliveira2013]; it also clearly differs from the PDFs in Fig.  8.
@@
We consider now the tails of the PDFs in some further detail. As noted from the values of skewness and kurtosis, these tails are asymmetric and non-Gaussian. For many random systems in their disorder-dominated phases, they are frequently seen to follow exponential functions as [Monthus2008]
@@
where $c_{\pm}$ are constants and $\eta_{\pm}$ are characteristic, so-called tail exponents [Monthus2008, Halpin15]. We have fitted the tails of the computed PDF to functions with the form given by Eq. (21) [dotted lines in Figs. 7(a) and 8(a)] to estimate the $\eta_{\pm}$ exponents; the results are collected in Table 3 as well. In relation to this, the tail exponents for the TW distribution have been characterized [majumdar2014, Halpin15]. For the $d=1$ KPZ universality class with time-dependent noise, the exponent of the left tail is $\eta_{-}=3$, while that of the so-called Airy tail (the right tail for TW-GOE in Fig. 7) is $\eta_{+}=3/2$ [kim1991, majumdar2014]. Moreover, they are related to each other for KPZ as $\eta_{-}=(d+1)\eta_{+}$, and to the growth exponent as $\eta_{+}=1/(1-\beta)$ [majumdar2014, Halpin15]. These relations do not hold for the $\eta_{\pm}$ and $\beta$ values collected in Tables 1, 2, and 3, although the fits are reasonable and such that $\eta_{-}>\eta_{+}>1$ for $d=1$ and 2, as implied by the KPZ formulae. Violations of the $\eta_{+}=1/(1-\beta)$ relation have been also reported, e.g., for spreading of precursor films in band geometry [Marcos22] or for synchronized oscillator lattices [Gutierrez2023] (for both of which the PDF is TW-GOE although the kinetic roughening exponents are not 1D KPZ), or for moving fronts associated with the contact process [Barreales2023].
@@ VI Discussion
1 DPD [Amaral95] 0.58(7) $1.01(2)$ $0.63(1)$ $0.63(1)$
DPD [Tang92, Buldyrev92] $\approx 0.636$ $\approx 1$ $\approx 0.633$ $\approx 0.633$
DPD [Sneppen92] $\approx 0.63$
qKPZ (eq.) [Lee05] 0.616(9) 0.360(1) $\approx 0.978$ 0.647(2) 0.633(8)
qKPZ (DM) [Song07] $\approx 0.672$ $\approx 0.377$ $\approx 1.02$ $\approx 0.617$ $\approx 0.627$
aDEP [Mukerjee2023] 0.635(6)
TL92 [Mukerjee2023] 1.10(2) 0.636(8)
qKPZ (eq.) [Mukerjee2023] 0.64(2)
@@
2 DPD [Amaral95] 0.8(2) 1.15(5) 0.41(5) 0.48(3)
DPD [Buldyrev93] $\approx 1.16$ $\approx 0.41$ $\approx 0.48$
aDEP [Mukerjee2023] 0.48(2)
TL92 [Mukerjee2023] 0.47(3)
qKPZ (eq., $F=F_{c}$) [Wu24] 1.03(2) 0.64(2) 1.32(3) 0.38(5) 0.50(3)
qKPZ (eq., $\lambda>\lambda_{c}$) [Wu24] 0.83(2) 0.58(2) 1.35(5) 0.41(1) 0.51(1)
@@
Table 4 collects our values for the scaling exponents relevant to the kinetic roughening behavior of the qKPZ universality class at depinning, for the two physical dimensions. The table allows for direct comparison with values previously reported in the literature from simulations of several different models in the same universality class —specifically, the discrete DPD model [Amaral95], the anisotropic depinning (aDEP) model [Rosso2001, Mukerjee2023], a cellular automaton proposed in Ref. [Tang1992] (TL92) and recently reconsidered [Mukerjee2023], as well as direct simulations of the qKPZ equation [Lee05, Wu24]—, which provide both well-known results and recent revisions thereof. At first sight, our results seem fairly consistent (albeit with reduced statistical errors) with those of the DPD model, and also with additional estimates; actually, the different sets of exponents are compatible with each other.
@@
In particular, our calculated values for $\alpha$ agree reasonably well with available experimental data for fluid imbibition in paper, which yield $\alpha_{\rm exp}=0.63(4)$ and $0.52(4)$ at depinning for $d=1$ and $d=2$, respectively [Amaral95]. More recent experiments obtain $\alpha_{\rm exp}=0.66\pm 0.04$ and $\beta_{\rm exp}=0.61\pm 0.05$ for $d=1$ reaction fronts in disordered flows [Atis2015].
@@
As we have already mentioned, one of the key features of our work is that we have been able to compute the entire set of scaling exponents directly from simulations, i.e., without employing any scaling relation a priori. This approach contrasts with that followed by previous authors. In their model for DPD, Amaral et al. [Amaral95] computed directly $\beta$ as reported in Table 4 (which are virtually identical to ours), but used scaling arguments for the longitudinal and transverse correlation lengths to estimate $\alpha$ for $d\leq 6$. For $z$ and $\theta$, they use again scaling arguments (see, for instance, Eqs. (3.9) and (3.13) in Ref. [Amaral95]), which may be the reason why their values for $\alpha$ and $z$ deviate more (by up to $\approx-3\%$ for $d=2$) from ours. This deviation is even larger for $\theta$ in $d=1$, for which their estimation is roughly 15% smaller than our calculation. On the other hand, Song and Kim [Song07] used an automaton model in $d=1$ to find $\beta\approx 0.617$ and $\alpha\approx 0.627$, and estimated $z=\alpha/\beta\approx 1.02$, about 2% higher than our calculated value. Similarly, Lee and Kim [Lee05] compute $\beta=0.647(2)$ and $\alpha=0.633(8)$ for the qKPZ equation in $d=1$ and estimate $z=\alpha/\beta\approx 0.978$, which is around 4% smaller than our value, the difference arising mostly from our dissimilar $\alpha$ values. Note that, according to our simulations, the $\alpha=\beta z$ and $\delta+\beta=1$ scaling relations do not hold exactly in the DM, especially for $d=2$. Given that scaling relations tend to be more accurately satisfied as $L$ increases, we attribute the small deviations found to residual finite-size effects.
@@
Some further discrepancies may be found in Table 4. In Ref. [Mukerjee2023], $z=1.10(2)$ was reported for the TL92 model, which is at 4.4 standard deviations away from our estimate and 5 standard deviations from the $z=1$ value conjectured in Ref. [Amaral95]. We remark that we have obtained our estimate for $z$ from an analysis of the correlation length extracted from the $C_{2}(\boldsymbol{r},t)$ correlation function; this analysis yields a value near $z=1$ for all the simulated lattice sizes, see Table 1. Besides, our final value, $z=1.016(7)$, is within two standard deviation of the $z=1$ conjectured DPD value. Obviously, we cannot prove the latter using numerical simulations; we can only disprove it, with a given statistical confidence level.
@@
On the other hand, Wu et al. [Wu24] computed the scaling exponents for the 2D model from the numerical simulation of the qKPZ equation. These authors address two different conditions, namely $F=F_{c}$ for $\lambda=5$ and $F=0$ at $\lambda=\lambda_{c}=2.865$, which may seem inconsistent with the conditions used in our present work. In this respect, note that a pinning-depinning transition can arise for the qKPZ equation with $F=0$, provided $\lambda>\lambda_{c}$ [Ramasco01, Wu24]. In this case, the system behaves as if it were subject to an effective driving force [Ramasco01], so that our data and those from Wu et al. can be properly compared. Actually, Wu et al. [Wu24] themselves suggest that the exponents they obtain for $\lambda=\lambda_{c}$ are “between those of qEW and qKPZ universality classes to some extent, and belong to the DPD class”. Considering both our present results and previous ones (see Table 4), the $\lambda=\lambda_{c}$ case of Ref. [Wu24] agrees well with the established values (with their $z$ exponent value being three standard deviations off our estimate). By contrast, their $F=F_{c}$ case shows noticeable discrepancies, most prominently in the $\delta$ exponent and, to a lesser degree, in the $z$ exponent.
@@
Finally, regarding the PDF of the front fluctuations, we remark that, to our knowledge, no previous analytical or numerical results exist for the qKPZ universality class at depinning. Our results can be naturally compared with those obtained for the standard KPZ universality class (with time-dependent noise), specifically with the TW-GOE distribution for $d=1$ [Kriecherbauer10, Halpin15, Takeuchi18] or, in the case of systems with absorbing states, with the PDF obtained [Barreales2023] for the dynamics of fronts associated with the contact process [Dickman2000]. In both cases the shape of the distributions and the values of skewness and kurtosis differ unambiguously from those we obtain for the qKPZ universality class, in the two space dimensions we have investigated. We expect that the PDF we are presently reporting can also be found at least in other models of the qKPZ universality class at depinning. For the sake of accessibility, we provide our numerical distributions $P(\chi)$ for $d=1$ and 2 in an open-access repository [datasetchi].

### v2

@@ III.3 Correlation function
Another check of the goodness of our results also comes from Eq. (18), which can be rewritten as $C_{2}(\boldsymbol{r},t)r^{-2\alpha}=g[\boldsymbol{r}/\xi(t)]$. Consequently, a plot of $C_{2}(\boldsymbol{r},t)r^{-2\alpha}$ vs. $r/\xi(t)$ should yield the universal (i.e., time-independent) scaling function $g(u)$ with asymptotic behaviors as indicated in Eq. (18). Figure 6(a) shows such a data collapse for $d=1$, for the selected times and different lattice sizes, in a doubly-logarithmic plot, in good agreement with the theoretical expectations. Note further that the $C_{2}(\boldsymbol{r},t)$ curves for different times share the same slope at small arguments; as a consequence, there is a range of $r$ where these curves partially overlap for different times. This indicates that the same scaling ansatz holds throughout the lattice, irrespective of the value of $r/\xi(t)$ and, in turn, that a single roughness exponent $\alpha$ suffices to describe the spatial properties of correlations in the system, as indeed expected within the FV dynamic scaling ansatz [5, 6, 52] (see also Fig. 10(a) in Appendix B).
@@
Following the same methodology used in Sec. III.1, for our simulations in $d=2$ we obtain $\theta=0.803(1)$ and $F_{c}=1.703(1)$ after fitting the $L=1024$ data to Eq. (12), see Fig. 1(b). To our knowledge, this result is new, as we are not aware of previous reports in the literature of the value of $F_{c}$ in $d=2$ for the model we are presently studying.
@@
No evidences of anomalous scaling has been observed, such as, for example, the characteristic anomalous upward shift of the $C_{2}(\boldsymbol{r},t)$ curves for increasing times [57, 58, 59]. This underscores the consistency of the FV dynamic scaling ansatz for the 2D model as well and contrasts with the non-FV, anomalous scaling observed for other systems with absorbing states, see, e.g., Ref. [49, 50] and others therein.
@@ V Statistics of fronts fluctuations
Besides the scaling exponents, the shape of the PDF for fluctuations of the front, defined according to Eq. (20) above, contributes substantially to the unambiguous identification of non-equilibrium universality classes, as has been demonstrated in several contexts [8, 9, 10, 26, 52, 27, 28, 50]. In particular, the skewness and excess kurtosis have been recently computed in a system closely related to the one we are addressing, namely, the quenched version of the Edwards-Wilkinson equation [60, 61].
Figure 7(a) shows the PDF for the front fluctuations computed for the 1D model at different times within the growth regime, at fixed $L=10^{6}$. This plot evidences that, first, once suitably rescaled, the distribution of fluctuations indeed becomes asymptotically independent of time and system size. Second, a markedly asymmetric shape is observed, with a sharp edge for negative $\chi$ and a more regular tail for positive $\chi$. As a reference, we have plotted in Fig. 7 the Tracy-Widom distributions for the largest eigenvalue of Hermitian random matrices in the Gaussian orthogonal (TW-GOE) and unitary (TW-GUE) ensembles, as well as the familiar Gaussian distribution. The first two are known [8, 9, 10] to describe the fluctuation PDF in the standard 1D KPZ universality class, in band geometry subject to PBC and for an overall circular symmetry of the front, respectively, whereas the Gaussian distribution holds for linear models of kinetically rough surfaces [5, 62], all of them in the case of time-dependent (rather than quenched) noise.
@@
Figure 7: PDF of the front fluctuations for the 1D model. (a) PDF computed for $L=10^{6}$ at two different times within the growth regime, see legend. The dotted lines correspond to the best fits of the tails of the PDF to Eq. (21), with the values for the tail exponents $\eta_{\pm}$ given in Table 3. (b) PDF computed at $t=50000$ for different system sizes, see legend. In both panels, the solid lines show the TW-GOE, TW-GUE, and Gaussian distributions as references.
Figure 8: PDF of the front fluctuations for the 2D model. (a) PDF computed for $L=512$ at two different times within the growth regime, see legend. The dotted lines correspond to the best fits of the tails of the PDF to Eq. (21), with the values for the tail exponents $\eta_{\pm}$ given in Table 3. (b) PDF computed at $t=600$ for different system sizes, see legend. In both panels and as references, the solid lines show the Gaussian distribution and the Gumbel’s first asymptotic form with parameter $m=6$ that describes quite accurately the fluctuation PDF for the 2D KPZ universality class [63] (green).
@@
Similar qualitative features are found for $d=2$, as shown in Fig. 8, where the PDFs exhibit quite different shapes, however. This is evidenced in Fig. 9, where $P(\chi)$ is plotted as a function of $\chi$ for both dimensionalities, with the data being available at the Zenodo open access repository [64]. Note that the PDF is known not to be in the TW family for the 2D KPZ universality class, despite being strongly non-Gaussian [65, 66, 63].
For comparison purposes, we have included in Fig. 8 the Gaussian distribution together with the so-called Gumbel’s first asymptotic distribution, as parametrized by Olivera et al. [63] which accurately describes the fluctuation properties of the 2D KPZ equation (with thermal noise). Notice that the fluctuations computed for the 2D qKPZ equation clearly differs from that of the thermal KPZ and are not Gaussian.
@@
We consider now the tails of the PDFs in some further detail. As noted from the values of skewness and kurtosis, these tails are asymmetric and non-Gaussian. For many random systems in their disorder-dominated phases, they are frequently seen to follow exponential functions as [67]
@@
where $c_{\pm}$ are constants and $\eta_{\pm}$ are characteristic, so-called tail exponents [67, 9]. We have fitted the tails of the computed PDF to functions with the form given by Eq. (21) [dotted lines in Figs. 7(a) and 8(a)] to estimate the $\eta_{\pm}$ exponents; the results are collected in Table 3 as well. In relation to this, the tail exponents for the TW distribution have been characterized [68, 9]. For the $d=1$ KPZ universality class with time-dependent noise, the exponent of the left tail is $\eta_{-}=3$, while that of the so-called Airy tail (the right tail for TW-GOE in Fig. 7) is $\eta_{+}=3/2$ [69, 68]. Moreover, they are related to each other for KPZ as $\eta_{-}=(d+1)\eta_{+}$, and to the growth exponent as $\eta_{+}=1/(1-\beta)$ [68, 9]. These relations do not hold for the $\eta_{\pm}$ and $\beta$ values collected in Tables 1, 2, and 3, although the fits are reasonable and such that $\eta_{-}>\eta_{+}>1$ for $d=1$ and 2, as implied by the KPZ formulae. Violations of the $\eta_{+}=1/(1-\beta)$ relation have been also reported, e.g., for spreading of precursor films in band geometry [28] or for synchronized oscillator lattices [17] (for both of which the PDF is TW-GOE although the kinetic roughening exponents are not 1D KPZ), or for moving fronts associated with the contact process [50].
@@ VI Discussion
1 DPD [42] 0.58(7) $1.01(2)$ $0.63(1)$ $0.63(1)$
DPD [70, 71] $\approx 0.636$ $\approx 1$ $\approx 0.633$ $\approx 0.633$
DPD [72] $\approx 0.63$
qKPZ (eq.) [48] 0.616(9) 0.360(1) $\approx 0.978$ 0.647(2) 0.633(8)
qKPZ (DM) [47] $\approx 0.672$ $\approx 0.377$ $\approx 1.02$ $\approx 0.617$ $\approx 0.627$
aDEP [45] 0.635(6)
TL92 [45] 1.10(2) 0.636(8)
qKPZ (eq.) [45] 0.64(2)
@@
2 DPD [42] 0.8(2) 1.15(5) 0.41(5) 0.48(3)
DPD [73] $\approx 1.16$ $\approx 0.41$ $\approx 0.48$
aDEP [45] 0.48(2)
TL92 [45] 0.47(3)
qKPZ (eq., $F=F_{c}$) [74] 1.03(2) 0.64(2) 1.32(3) 0.38(5) 0.50(3)
qKPZ (eq., $\lambda>\lambda_{c}$) [74] 0.83(2) 0.58(2) 1.35(5) 0.41(1) 0.51(1)
@@
Table 4 collects our values for the scaling exponents relevant to the kinetic roughening behavior of the qKPZ universality class at depinning, for the two physical dimensions. The table allows for direct comparison with values previously reported in the literature from simulations of several different models in the same universality class —specifically, the discrete DPD model [42], the anisotropic depinning (aDEP) model [75, 45], a cellular automaton proposed in Ref. [76] (TL92) and recently reconsidered [45], as well as direct simulations of the qKPZ equation [48, 74]—, which provide both well-known results and recent revisions thereof. At first sight, our results seem fairly consistent (albeit with reduced statistical errors) with those of the DPD model, and also with additional estimates; actually, the different sets of exponents are compatible with each other.
@@
In particular, our calculated values for $\alpha$ agree reasonably well with available experimental data for fluid imbibition in paper, which yield $\alpha_{\rm exp}=0.63(4)$ and $0.52(4)$ at depinning for $d=1$ and $d=2$, respectively [42]. More recent experiments obtain $\alpha_{\rm exp}=0.66\pm 0.04$ and $\beta_{\rm exp}=0.61\pm 0.05$ for $d=1$ reaction fronts in disordered flows [35].
@@
As we have already mentioned, one of the key features of our work is that we have been able to compute the entire set of scaling exponents directly from simulations, i.e., without employing any scaling relation a priori. This approach contrasts with that followed by previous authors. In their model for DPD, Amaral et al. [42] computed directly $\beta$ as reported in Table 4 (which are virtually identical to ours), but used scaling arguments for the longitudinal and transverse correlation lengths to estimate $\alpha$ for $d\leq 6$. For $z$ and $\theta$, they use again scaling arguments (see, for instance, Eqs. (3.9) and (3.13) in Ref. [42]), which may be the reason why their values for $\alpha$ and $z$ deviate more (by up to $\approx-3\%$ for $d=2$) from ours. This deviation is even larger for $\theta$ in $d=1$, for which their estimation is roughly 15% smaller than our calculation. On the other hand, Song and Kim [47] used an automaton model in $d=1$ to find $\beta\approx 0.617$ and $\alpha\approx 0.627$, and estimated $z=\alpha/\beta\approx 1.02$, about 2% higher than our calculated value. Similarly, Lee and Kim [48] compute $\beta=0.647(2)$ and $\alpha=0.633(8)$ for the qKPZ equation in $d=1$ and estimate $z=\alpha/\beta\approx 0.978$, which is around 4% smaller than our value, the difference arising mostly from our dissimilar $\alpha$ values. Note that, according to our simulations, the $\alpha=\beta z$ and $\delta+\beta=1$ scaling relations do not hold exactly in the DM, especially for $d=2$. Given that scaling relations tend to be more accurately satisfied as $L$ increases, we attribute the small deviations found to residual finite-size effects.
@@
Some further discrepancies may be found in Table 4. In Ref. [45], $z=1.10(2)$ was reported for the TL92 model, which is at 4.4 standard deviations away from our estimate and 5 standard deviations from the $z=1$ value conjectured in Ref. [42]. We remark that we have obtained our estimate for $z$ from an analysis of the correlation length extracted from the $C_{2}(\boldsymbol{r},t)$ correlation function; this analysis yields a value near $z=1$ for all the simulated lattice sizes, see Table 1. Besides, our final value, $z=1.016(7)$, is within two standard deviation of the $z=1$ conjectured DPD value. Obviously, we cannot prove the latter using numerical simulations; we can only disprove it, with a given statistical confidence level.
@@
On the other hand, Wu et al. [74] computed the scaling exponents for the 2D model from the numerical simulation of the qKPZ equation. These authors address two different conditions, namely $F=F_{c}$ for $\lambda=5$ and $F=0$ at $\lambda=\lambda_{c}=2.865$, which may seem inconsistent with the conditions used in our present work. In this respect, note that a pinning-depinning transition can arise for the qKPZ equation with $F=0$, provided $\lambda>\lambda_{c}$ [39, 74]. In this case, the system behaves as if it were subject to an effective driving force [39], so that our data and those from Wu et al. can be properly compared. Actually, Wu et al. [74] themselves suggest that the exponents they obtain for $\lambda=\lambda_{c}$ are “between those of qEW and qKPZ universality classes to some extent, and belong to the DPD class”. Considering both our present results and previous ones (see Table 4), the $\lambda=\lambda_{c}$ case of Ref. [74] agrees well with the established values (with their $z$ exponent value being three standard deviations off our estimate). By contrast, their $F=F_{c}$ case shows noticeable discrepancies, most prominently in the $\delta$ exponent and, to a lesser degree, in the $z$ exponent.
@@
Finally, regarding the PDF of the front fluctuations, we remark that, to our knowledge, no previous analytical or numerical results exist for the qKPZ universality class at depinning. Our results can be naturally compared with those obtained for the standard KPZ universality class (with time-dependent noise), specifically with the TW-GOE distribution for $d=1$ [8, 9, 10] or, in the case of systems with absorbing states, with the PDF obtained [50] for the dynamics of fronts associated with the contact process [49]. In both cases the shape of the distributions and the values of skewness and kurtosis differ unambiguously from those we obtain for the qKPZ universality class, in the two space dimensions we have investigated. We expect that the PDF we are presently reporting can also be found at least in other models of the qKPZ universality class at depinning. For the sake of accessibility, we provide our numerical distributions $P(\chi)$ for $d=1$ and 2 in an open-access repository [64].